OpenStudy (anonymous):

Problem set 3, question 2A-1 says: Find the linearization of sqrt(a + bx) at 0, by using (2), and also by using the basic approximation formulas. (Here a and b are constants; assume a > 0. Do not confuse this a with the one in (2), which has the value 0.) 1. what does "(2)" refer to? 2. one of the solutions given is: sqrt(a)sqrt(1 + bx/a) . How is this calculated? Thanks

OpenStudy (anonymous):

sqrt(a+bx) = sqrt (a(1+bx/a)) = sqrt(a)sqrt(1+bx/a) I guess (2) refers to using the generic formula...

OpenStudy (anonymous):

Thanks. That answers it.